10. Use the table below. The Bureau of Labor Statistics reports on a variety of employment statistics. College Enrollment anIs employment independent of education for women in the civilian labor force? (Show this mathematically) (2 points)
Answer #1

The probability of a woman employed = 4985/(4985+720) = 0.874

The probability of a woman employed = (1352+921)/(4985+720) = 0.398

No, employment is not independent of education for women in the civilian labor force. For example:

The probability of women employed who is in the civilian labor force who has some college degree = 1352/(4985+720) = 0.237

Please give me a thumbs-up if this helps you out. Thank you!